# Grid-Based Continuous Normal Representation for Anomaly Detection

### v1

@@ 4.2 Effectiveness of Continuous Feature Space
Cable 60.0/86.2 70.9/92.7 78.2/89.2 57.8/91.8 85.0/95.0 99.7/99.4 95.2/97.6 99.0/- 99.5±0.15/99.8
Capsule 57.6/86.1 73.4/91.3 68.3/80.5 65.3/98.5 95.5/96.3 90.9/97.8 86.9/85.3 95.4/- 96.9±0.37/97.9
Hazelnut 95.8/93.1 85.5/92.0 97.1/98.4 93.7/100 87.1/99.9 100/100 99.8/99.9 100/- 99.9±0.09/100
Metal Nut 62.7/82.0 88.0/98.7 64.9/73.6 72.8/98.7 99.4/100 99.9/100 99.2/99.0 99.9/- 99.9±0.08/99.9
Pill 56.1/87.9 68.8/93.3 79.7/82.7 82.2/98.9 52.6/96.6 96.9/96.0 93.7/88.3 95.8/- 97.8±0.37/98.7
Screw 66.9/54.9 56.9/85.8 75.6/83.3 92/93.9 97.3/97.0 90.1/97.0 87.5/91.9 95.6/- 97.5±0.43/99.0
Toothbrush 57.8/95.3 95.3/96.1 75.3/92.2 90.6/100 99.4/99.5 100/99.7 94.2/95.0 93.6/- 99.3±0.13/97.2
Transistor 61.0/81.8 86.6/97.4 73.4/85.6 74.8/93.1 92.4/96.7 99.7/100 99.8/100 99.7/- 99.9±0.09/99.9
Zipper 78.6/91.9 79.7/90.3 87.4/93.2 98.8/100 99.6/98.5 94.7/99.5 95.8/96.7 97.9/- 99.2±0.15/99.2
Carpet 86.6/91.6 93.8/99.8 69.8/79.3 98.0/97.0 97.1/98.9 97.1/98.7 99.8/99.9 99.9/- 100±0.03/100
@@
Class US [4] PaDiM [8] MKD [32] DRAEM [41] RD4AD [9] PatchCore [30] UniAD [40] HVQ-T [23] GRAD (Ours)
Bottle 67.9/97.8 96.1/98.2 91.8/96.3 87.6/99.1 97.7/98.7 98.4/98.6 98.1/98.1 98.3/- 98.4±0.11/98.5
Cable 78.3/91.9 81.0/96.7 89.3/82.4 71.3/94.7 83.1/97.4 96.7/98.5 97.3/96.8 98.1/- 98.3±0.16/98.3
Capsule 85.5/96.8 96.9/98.6 88.3/95.9 50.5/94.3 98.5/98.7 94.8/98.9 98.5/97.9 98.8/- 98.6±0.02/98.4
Hazelnut 93.7/98.2 96.3/98.1 91.2/94.6 96.9/99.7 98.7/98.9 98.6/98.7 98.1/98.8 98.8/- 98.7±0.19/98.7
Metal Nut 76.6/97.2 84.8/97.3 64.2/86.4 62.2/99.5 94.1/97.3 98.3/98.4 94.8/95.7 96.3/- 97.6±0.21/97.8
Pill 80.3/96.5 87.7/95.7 69.7/89.6 94.4/97.6 96.5/98.2 97.3/97.6 95.0/95.1 97.1/- 97.9±0.10/97.9
Screw 90.8/97.4 94.1/98.4 92.1/96.0 95.5/97.6 99.4/99.6 98.0/99.4 98.3/97.4 98.9/- 99.2±0.06/99.1
Toothbrush 86.9/97.9 95.6/98.8 88.9/96.1 97.7/98.1 99.0/99.1 98.4/98.7 98.4/97.8 98.6/- 98.8±0.03/94.1
Transistor 68.3/73.7 92.3/97.6 71.7/76.5 64.5/90.9 86.4/92.5 94.9/96.4 97.9/98.7 97.9/- 98.3±0.25/98.7
Zipper 84.2/95.6 94.8/98.4 86.1/93.9 98.3/98.8 98.1/98.2 95.8/98.9 96.8/96.0 97.5/- 97.9±0.07/97.7
Carpet 88.7/93.5 97.6/99.0 95.5/95.6 98.6/95.5 98.8/98.9 98.9/99.1 98.5/98.0 98.7/- 98.7±0.02/98.8
@@
Leather 95.4/97.8 84.8/99.0 96.7/98.1 97.3/98.6 99.4/99.4 99.3/99.3 98.8/98.3 98.8/- 98.7±0.20/99.2
Tile 82.7/92.5 80.5/94.1 85.3/82.8 98.0/99.2 95.6/95.6 95.9/95.9 91.8/91.8 92.2/- 94.2±0.20/94.1
Wood 83.3/92.1 89.1/94.1 80.5/84.8 96.0/96.4 96.0/95.3 94.4/95.1 93.2/93.4 92.4/- 94.0±0.10/94.6
@@
Sampled features from the grids (normal vs. abnormal). In addition, we visualize the sampled features from the local and global grids based on the learned coordinates. Fig. 3 shows that the sampled features (from the local and global grids) with near coordinates share similar characteristics whether the input image is normal or not. Furthermore, the fused normal representations of both normal and abnormal inputs are reconstructed into normal images. This result demonstrates that the continuous feature space efficiently tackles the two major challenges in discrete feature space: weak generalization and IS. The first challenge is resolved by preserving the fine-grained details of the bottle, while the second is addressed by reconstructing the anomalous region that is never encountered during training back to its corresponding normal state.
Decision boundary of multiple classes. Fig. 4 describes the coordinate distribution using a model trained with global representation. The images from each class form clusters in the continuous feature space, effectively modeling the decision boundaries between classes. This implies that the continuous feature space can represent well-defined structural features. Furthermore, the reconstructions of the sampled features at the decision boundary show combined characteristics of near classes, demonstrating the high granularity of continuous features. Leveraging the advantages of the continuous feature space, GRAD can model correct decision boundaries in complex multi-class distributions with compact representations, leading to high performance in a unified setting.
Figure 5: Qualitative results of GRAD on MVTec AD. Each row of the figure represents anomaly images, corresponding ground truths, results from UniAD, and our results.
@@ 4.3 Anomaly Detection and Localization
We evaluate GRAD in comparison with recent state-of-the-art methods in both unified and separate settings, focusing on detection (Tab. 2) and localization (Tab. 3) performance on MVTec AD. In a unified setting, the methods not specifically designed for multiple classes exhibit a significant performance drop compared to their performance in a separate setting. In contrast, UniAD, HVQ-Trans, and GRAD maintain the performances of their separate models in the unified setting. Among these, GRAD notably outperforms UniAD and HVQ-Trans, achieving state-of-the-art performance in the unified setting. Specifically, GRAD successfully reduces the error rate of HVQ-Trans from 2.0% to 0.7%, bringing a total error reduction of 65.0%. For detection, the unified GRAD even outperforms separate models of PatchCore, which is the previous state-of-the-art in single-class AD. Similarly, for localization, GRAD achieves the best performance in the unified setting and matches PatchCore in a separate setting. Fig. 5 showcases the qualitative results of UniAD and GRAD in the unified setting, highlighting GRAD’s superior prediction quality with fewer noisy areas. We additionally evaluate GRAD on VisA, which is a more challenging dataset. As shown in Tab. 4, GRAD outperforms other state-of-the-art methods for detection.
@@
UniAD [40] PatchCore [30] OmniAL [42] GRAD (ours) UniAD [40] PatchCore [30] OmniAL [42] GRAD (ours)
Complex Structure PCB1 94.8/90.2 97.6/98.5 77.7/96.6 98.0/95.3 99.3/99.2 99.7/99.8 97.6/98.7 99.5/99.3
PCB2 92.5/84.2 96.7/97.2 81.0/99.4 90.5/93.3 97.6/96.5 98.0/98.7 93.9/83.2 97.3/96.6
PCB3 86.6/90.7 97.3/98.5 88.1/96.9 96.2/96.5 98.1/98.0 99.3/99.4 94.7/98.4 98.7/98.5
PCB4 99.3/97.4 99.7/99.7 95.3/97.4 99.1/98.1 97.6/97.2 97.7/98.2 97.1/98.5 98.3/98.0
Multiple Instances Macaroni1 90.4/90.2 94.7/97.4 92.6/96.9 93.5/95.2 99.1/99.0 99.0/99.7 98.6/98.9 99.0/99.0
Macaroni2 82.8/77.4 78.6/76.7 75.2/89.9 95.9/88.8 97.7/97.4 96.1/98.6 97.9/99.1 98.8/98.8
Capsules 70.7/80.3 75.0/76.3 90.6/87.9 87.3/93.5 98.1/98.5 99.1/99.2 99.4/98.6 99.4/99.6
Candle 97.0/90.2 94.7/99.4 86.8/85.1 95.9/96.9 99.1/99.0 98.3/99.3 95.8/90.5 99.3/99.2
Single Instance Cashew 93.8/92.9 97.3/97.8 88.6/97.1 93.6/96.4 98.9/99.2 98.1/98.7 95.0/98.9 97.8/97.5
Chewinggum 99.3/98.3 98.5/98.8 96.4/94.9 94.6/99.6 99.1/98.5 98.9/98.9 99.0/98.7 98.3/98.6
Fryum 88.8/84.4 95.4/96.0 94.6/97.0 99.5/93.2 97.7/96.7 89.8/92.4 92.1/89.3 96.7/96.1
Pipe fryum 97.0/91.8 99.2/99.8 86.1/91.4 98.0/98.8 99.3/99.3 97.5/98.9 98.2/99.1 99.4/98.9
Mean 91.1/89.0 93.7/94.7 87.8/94.2 95.2/95.5 98.5/98.2 97.5/98.5 96.6/96.0 98.5/98.3
@@
Table 5: Ablation studies on the proposed method in the unified setting using MVTec AD.
4.4 Ablation Study
We conducted an ablation study on GRAD to assess the impact of individual proposals. Tab. 5 shows that our key contribution is the normal representation from both local and global contexts, which independently yields comparable performance. Notably, a model with only local representation outperforms UniAD and matches HVQ-Trans. The integration of both representations achieves improved performance, with additional gains from feature refinement and coordinate jittering.
@@ 5 Conclusion
In this work, we have proposed a novel anomaly detection architecture, GRAD, which represents normal features in the continuous space, unlike prior approaches limited to discrete space. GRAD successfully represents local as well as global features in the continuous space while overcoming the limitations of existing methods, such as weak generalization, identity shortcut, computational complexity, and parameter efficiency. Through extensive experiments, we have demonstrated the effectiveness of GRAD qualitatively and quantitatively. With feature refinement and coordinate jittering, GRAD achieves state-of-the-art performance by a significant margin.
@@ Appendix 0.A Implementation Details
We implemented GRAD in the PyTorch framework, and we used the NVIDIA A5000 GPU for all evaluations.
Training. We trained our models for 100 epochs, thrice with different seeds (0,1,2), with a batch size of 64. The learning rate is initially set to $1\times 10^{-3}$ and $1\times 10^{-1}$ for neural networks and grids, respectively, which are reduced by a factor of 0.1 twice at the 40 and 80 epochs. We used AdamW optimizer [22] with weight decay $1\times 10^{-2}$.
Model configurations. To create the input feature $x$, we utilized the third and fourth stage feature maps of EfficientNet-b4 [36], which are resized to $28\times 28$ and then concatenated. We configured both the dimensions of the local ($C_{l}$) and global coordinates ($C_{g}$) to 2. Additionally, we set the grid resolution of the local ($R_{l}$) and global ($R_{g}$) representations to 8 and 4, respectively. In terms of feature refinement, we set the parameters $\lambda_{1}$, $\lambda_{2}$, and $k$ to be 0.3, 0.7, and 10. For coordinate jittering, we applied Gaussian noise $N(0,1)$ scaled by a factor of 0.05 to 50% of pixel-wise features. We designed the convolutional network $\psi$ with 7 basic residual blocks [14] that maintain the number of channels and the spatial resolution. Unless otherwise noted, the model performances were measured in a unified setting.
@@ 0.B.1 Computational and Parameter Efficiency
We evaluated the model size and computational complexity of GRAD. As shown in Tab. 6, GRAD requires reasonable storage and computation resources. However, for a fair comparison with UniAD, we downsized the convolutional networks by half the channels while maintaining the grid configurations, coined as GRAD-S. GRAD-S exhibits superior performance with reduced memory and computation requirements, demonstrating both effectiveness and efficiency. Furthermore, GRAD has another strong point: fast training time enabled by grids. The outstanding performance is achieved only with $1/10$ training duration compared to UniAD.

### v3

@@ 4.2 Effectiveness of Continuous Memory Representation
Cable 60.0/86.2 70.9/92.7 78.2/89.2 57.8/91.8 85.0/95.0 99.7/99.4 95.2/97.6 99.0/- 99.1±0.34/99.7
Capsule 57.6/86.1 73.4/91.3 68.3/80.5 65.3/98.5 95.5/96.3 90.9/97.8 86.9/85.3 95.4/- 97.0±0.05/98.4
Hazelnut 95.8/93.1 85.5/92.0 97.1/98.4 93.7/100 87.1/99.9 100/100 99.8/99.9 100/- 100±0.06/100
Metal Nut 62.7/82.0 88.0/98.7 64.9/73.6 72.8/98.7 99.4/100 99.9/100 99.2/99.0 99.9/- 100±0.00/100
Pill 56.1/87.9 68.8/93.3 79.7/82.7 82.2/98.9 52.6/96.6 96.9/96.0 93.7/88.3 95.8/- 98.6±0.36/98.7
Screw 66.9/54.9 56.9/85.8 75.6/83.3 92/93.9 97.3/97.0 90.1/97.0 87.5/91.9 95.6/- 97.6±0.33/98.6
Toothbrush 57.8/95.3 95.3/96.1 75.3/92.2 90.6/100 99.4/99.5 100/99.7 94.2/95.0 93.6/- 99.2±0.73/96.1
Transistor 61.0/81.8 86.6/97.4 73.4/85.6 74.8/93.1 92.4/96.7 99.7/100 99.8/100 99.7/- 99.8±0.18/99.9
Zipper 78.6/91.9 79.7/90.3 87.4/93.2 98.8/100 99.6/98.5 94.7/99.5 95.8/96.7 97.9/- 99.2±0.13/99.6
Carpet 86.6/91.6 93.8/99.8 69.8/79.3 98.0/97.0 97.1/98.9 97.1/98.7 99.8/99.9 99.9/- 99.9±0.05/100
@@
Class US [4] PaDiM [8] MKD [30] DRAEM [38] RD4AD [9] PatchCore [28] UniAD [37] HVQ-T [21] CRAD (Ours)
Bottle 67.9/97.8 96.1/98.2 91.8/96.3 87.6/99.1 97.7/98.7 98.4/98.6 98.1/98.1 98.3/- 98.2±0.10/98.6
Cable 78.3/91.9 81.0/96.7 89.3/82.4 71.3/94.7 83.1/97.4 96.7/98.5 97.3/96.8 98.1/- 98.4±0.17/98.3
Capsule 85.5/96.8 96.9/98.6 88.3/95.9 50.5/94.3 98.5/98.7 94.8/98.9 98.5/97.9 98.8/- 98.7±0.08/98.6
Hazelnut 93.7/98.2 96.3/98.1 91.2/94.6 96.9/99.7 98.7/98.9 98.6/98.7 98.1/98.8 98.8/- 98.5±0.17/98.9
Metal Nut 76.6/97.2 84.8/97.3 64.2/86.4 62.2/99.5 94.1/97.3 98.3/98.4 94.8/95.7 96.3/- 97.5±0.36/97.3
Pill 80.3/96.5 87.7/95.7 69.7/89.6 94.4/97.6 96.5/98.2 97.3/97.6 95.0/95.1 97.1/- 98.2±0.03/98.0
Screw 90.8/97.4 94.1/98.4 92.1/96.0 95.5/97.6 99.4/99.6 98.0/99.4 98.3/97.4 98.9/- 99.3±0.04/99.2
Toothbrush 86.9/97.9 95.6/98.8 88.9/96.1 97.7/98.1 99.0/99.1 98.4/98.7 98.4/97.8 98.6/- 98.8±0.04/98.7
Transistor 68.3/73.7 92.3/97.6 71.7/76.5 64.5/90.9 86.4/92.5 94.9/96.4 97.9/98.7 97.9/- 98.1±0.14/98.3
Zipper 84.2/95.6 94.8/98.4 86.1/93.9 98.3/98.8 98.1/98.2 95.8/98.9 96.8/96.0 97.5/- 97.8±0.06/97.9
Carpet 88.7/93.5 97.6/99.0 95.5/95.6 98.6/95.5 98.8/98.9 98.9/99.1 98.5/98.0 98.7/- 98.6±0.06/98.7
@@
Leather 95.4/97.8 84.8/99.0 96.7/98.1 97.3/98.6 99.4/99.4 99.3/99.3 98.8/98.3 98.8/- 98.9±0.07/99.1
Tile 82.7/92.5 80.5/94.1 85.3/82.8 98.0/99.2 95.6/95.6 95.9/95.9 91.8/91.8 92.2/- 94.4±0.16/94.6
Wood 83.3/92.1 89.1/94.1 80.5/84.8 96.0/96.4 96.0/95.3 94.4/95.1 93.2/93.4 92.4/- 93.8±0.09/93.8
@@
Sampled features from the grids (normal vs. abnormal). In addition, we visualize the sampled features from the local and global grids based on the learned coordinates. Fig. 3 shows that the sampled features (from the local and global grids) with near coordinates share similar characteristics whether the input image is normal or not. Furthermore, the fused normal representations of both normal and abnormal inputs are reconstructed into normal images. Specifically, CRAD preserves the fine-grained details of the bottle (normal region), while it reconstructs the corresponding normal state of the anomalous region that has never been encountered during training. This result demonstrates that the continuous feature space efficiently tackles the two major challenges in discrete feature space: weak generalization and IS.
Decision boundary of multiple classes. Fig. 4 describes the coordinate distribution using a model trained with global representation. The images from each class form clusters in the continuous memory space, effectively modeling the decision boundaries between classes. This implies that the continuous memory can represent well-defined structural features. Furthermore, the reconstructions of the sampled features at the decision boundary show combined characteristics of near classes, demonstrating the high granularity of continuous features. Leveraging the advantages of the continuous feature memory, CRAD can model correct decision boundaries in complex multi-class distributions with compact representations, leading to high performance in a unified setting.
Figure 5: Qualitative results of CRAD on MVTec AD. Each row of the figure represents anomaly images, corresponding ground truths, results from UniAD, and our results.
@@ 4.3 Anomaly Detection and Localization
We evaluate CRAD in comparison with recent state-of-the-art methods in both unified and separate settings, focusing on detection (Tab. 2) and localization (Tab. 3) performance on MVTec AD. In the unified setting, the methods not specifically designed for multiple classes exhibit a significant performance drop compared to their performance in the separate setting. In contrast, UniAD, HVQ-Trans, and CRAD maintain the performances of their separate models in the unified setting. Among these, CRAD notably outperforms UniAD and HVQ-Trans, achieving state-of-the-art performance in the unified setting. Specifically, CRAD successfully reduces the error rate of HVQ-Trans from 2.0% to 0.7%, bringing a total error reduction of 65.0%. For detection, the unified CRAD even outperforms separate models of PatchCore, which is the previous state-of-the-art in single-class AD. Similarly, for localization, CRAD achieves the best performance in the unified setting and matches PatchCore in a separate setting. Fig. 5 showcases the qualitative results of UniAD and CRAD in the unified setting, highlighting CRAD’s superior prediction quality with fewer noisy areas. We additionally evaluate CRAD on VisA, which is a more challenging dataset. As shown in Tab. 4, CRAD outperforms other state-of-the-art methods for detection.
@@
UniAD [37] PatchCore [28] OmniAL [39] CRAD (ours) UniAD [37] PatchCore [28] OmniAL [39] CRAD (ours)
Complex Structure PCB1 94.8/90.2 97.6/98.5 77.7/96.6 96.8/95.4 99.3/99.2 99.7/99.8 97.6/98.7 99.5/99.5
PCB2 92.5/84.2 96.7/97.2 81.0/99.4 92.9/92.7 97.6/96.5 98.0/98.7 93.9/83.2 97.6/97.0
PCB3 86.6/90.7 97.3/98.5 88.1/96.9 95.2/96.1 98.1/98.0 99.3/99.4 94.7/98.4 98.7/98.6
PCB4 99.3/97.4 99.7/99.7 95.3/97.4 99.4/98.6 97.6/97.2 97.7/98.2 97.1/98.5 98.6/98.4
Multiple Instances Candle 97.0/90.2 94.7/99.4 86.8/85.1 96.3/96.6 99.1/99.0 98.3/99.3 95.8/90.5 99.2/99.2
Capsules 70.7/80.3 75.0/76.3 90.6/87.9 90.5/91.5 98.1/98.5 99.1/99.2 99.4/98.6 99.5/99.5
Macaroni1 90.4/90.2 94.7/97.4 92.6/96.9 96.6/96.0 99.1/99.0 99.0/99.7 98.6/98.9 99.1/99.1
Macaroni2 82.8/77.4 78.6/76.7 75.2/89.9 88.7/90.4 97.7/97.4 96.1/98.6 97.9/99.1 98.8/99.0
Single Instance Cashew 93.8/92.9 97.3/97.8 88.6/97.1 95.5/96.4 98.9/99.2 98.1/98.7 95.0/98.9 97.4/98.0
Chewinggum 99.3/98.3 98.5/98.8 96.4/94.9 99.5/98.9 99.1/98.5 98.9/98.9 99.0/98.7 98.3/98.4
Fryum 88.8/84.4 95.4/96.0 94.6/97.0 94.5/93.7 97.7/96.7 89.8/92.4 92.1/89.3 96.6/96.3
Pipe fryum 97.0/91.8 99.2/99.8 86.1/91.4 96.6/98.3 99.3/99.3 97.5/98.9 98.2/99.1 99.4/99.4
Mean 91.1/89.0 93.7/94.7 87.8/94.2 95.2/95.4 98.5/98.2 97.5/98.5 96.6/96.0 98.6/98.5
@@
4.4 Ablation Study
We conducted an ablation study on CRAD to assess the impact of individual proposals. Sec. 4.4 shows that our key contribution is the normal representation from both local and global contexts, which independently yields comparable performance. Notably, a model with only local representation outperforms UniAD and HVQ-
Table 5: Ablation studies in the unified setting using MVTec AD.
@@ 4.4 Ablation Study
Trans. The integration of both representations achieves improved performance, with additional gains from feature refinement and coordinate jittering.
@@ 5 Conclusion
In this work, we have proposed a novel anomaly detection architecture, CRAD, which represents normal features in the continuous memory, unlike prior approaches limited to discrete feature space. CRAD successfully represents local as well as global features in the continuous space while overcoming the limitations of existing methods, such as weak generalization, identity shortcut, and high computational/parameter complexity. Through extensive experiments, we have demonstrated the effectiveness of CRAD qualitatively and quantitatively. Although CRAD demonstrates its superior generalization capability compared to existing methods, we found a limitation that this cannot be the case with extremely limited data (i.e., 1- or zero-shot), more discussed in the supplementary materials. We believe that it can be further addressed and our work paves the way for future advancements in anomaly detection.
Acknowledgements
This work was supported in part by the Institute of Information and Communications Technology Planning and Evaluation (IITP) funded by the Korea Government (MSIT) under Grant RS-2021-II212068 (Artificial Intelligence Innovation Hub) and Grant RS-2022-II220688 (AI Platform to Fully Adapt and Reflect Privacy-Policy Changes); in part by the Culture, Sports, and Tourism R&D Program through the Korea Creative Content Agency funded by the Ministry of Culture, Sports and Tourism in 2024 under Grant RS-2024-00348469 (Research on neural watermark technology for copyright protection of generative AI 3D content); and in part by the SEMES-Sungkyunkwan University collaboration funded by SEMES.
@@ Appendix 0.A Implementation Details
We implemented CRAD in the PyTorch framework, and we used the NVIDIA A5000 GPU for all evaluations.
Training. We trained our models for 50 epochs, thrice with different seeds (0,1,2), with a batch size of 64. The learning rate is initially set to $1\times 10^{-3}$ and $1\times 10^{-1}$ for neural networks and grids, respectively, which are reduced by a factor of 0.1 once at the 40 epoch. We used AdamW optimizer [adamw] with weight decay $1\times 10^{-2}$.
Model configurations. To create the input feature $x$, we utilized the third and fourth stage feature maps of EfficientNet-b4 [effinet], which are resized to $28\times 28$ and then concatenated (channel $C=216$). We configured both the dimensions of the local ($C_{l}$) and global coordinates ($C_{g}$) to 2. Additionally, we set the grid resolution of the local ($R_{l}$) and global ($R_{g}$) representations to 8 and 4, respectively. In terms of feature refinement, we set the parameters $\lambda_{1}$, $\lambda_{2}$, and $k$ to be 0.3, 0.7, and 10. For coordinate jittering, we applied Gaussian noise $N(0,1)$ scaled by a factor of 0.05 to 50% of pixel-wise features. We designed the convolutional network $\psi$ with 11 MobileNetV3 [resnet] blocks that maintain the number of channels and the spatial resolution. Unless otherwise noted, the model performances were measured in a unified setting.
@@ 0.B.1 Computational and Parameter Efficiency
We evaluated the model size and computational complexity of CRAD. As shown in Tab. 6, CRAD requires reasonable storage and computation resources. However, for a fair comparison with UniAD, we downsized the convolutional networks while maintaining the grid configurations, coined as CRAD-S. CRAD-S exhibits superior performance with reduced memory and computation requirements, demonstrating both effectiveness and efficiency. Furthermore, CRAD has another strong point: fast training time enabled by grids. The outstanding performance is achieved only with $1/20$ training duration compared to UniAD.